### v4

@@ Abstract
This paper presents the design of a hardware accelerator for Transformers, optimized for on-device time-series forecasting in AIoT systems. It integrates integer-only quantization and Quantization-Aware Training with optimized hardware designs to realize 6-bit and 4-bit quantized Transformer models, which achieved precision comparable to 8-bit quantized models from related research. Utilizing a complete implementation on an embedded FPGA (Xilinx Spartan-7 XC7S15), we examine the feasibility of deploying Transformer models on embedded IoT devices. This includes a thorough analysis of achievable precision, resource utilization, timing, power, and energy consumption for on-device inference. Our results indicate that while sufficient performance can be attained, the optimization process is not trivial. For instance, reducing the quantization bitwidth does not consistently result in decreased latency or energy consumption, underscoring the necessity of systematically exploring various optimization combinations. Compared to an 8-bit quantized Transformer model in related studies, our 4-bit quantized Transformer model increases test loss by only 0.63%, operates up to 132.33$\times$ faster, and consumes 48.19$\times$ less energy.
Index Terms:
AIoT, Time-series Forecasting, Transformer, Integer-only Quantization, Embedded FPGAs
@@ I Introduction
  • We analyze the feasibility of our approach by deploying our generated accelerators on a Xilinx Spartan-7 XC7S15 FPGA. This FPGA is too small to deploy larger variants of our Transformer models, forcing us to compromise between precision and size, similar to a real IoT system. Our results show that we can deploy a 4-bit quantized Transformer model that increases test loss by only 0.63% but is 132.33$\times$ faster and consumes 48.19$\times$ less energy.
@@ IV-A Addition Component
$A^{3}=A^{1}+A^{2}\$ (7)
$A^{3}\approx S_{A^{1}}\cdot(A^{1}_{q}-Z_{A^{1}})+S_{A^{2}}\cdot(A^{2}_{q}-Z_{A^{2}})\$ (8)
$A^{3}_{q}\approx\left(\frac{S_{A^{1}}}{S_{A^{3}}}\cdot(A^{1}_{q}-Z_{A^{1}})+\frac{S_{A^{2}}}{S_{A^{3}}}\cdot(A^{2}_{q}-Z_{A^{2}})\right)+Z_{A^{3}}\$ (9)
@@ IV-C Softmax Component
$S_{E}=\frac{1}{((2^{(2b-1)}-1)-(-2^{(2b-1)}))/(n^{2}\cdot h)}\$ (12)
$Z_{E}=2^{(2b-1)}-\frac{1}{S_{E}}\$ (13)
$\text{DLUT}(\hat{X}_{q})=\text{clamp}(\text{round}(\frac{E}{S_{E}}),-2^{2b-1},2^{2b-1}-1)\$ (14)
$\text{NLUT}(\hat{X}_{q})=\text{clamp}(\text{round}(\frac{E}{S_{E}\cdot S_{A}}),-2^{3b-1},2^{3b-1}-1)\$ (15)
@@ IV-D Batch Normalization Component
$A_{q}(i,j)$ $\approx\frac{S_{\hat{\gamma_{j}}}\cdot S_{X}}{S_{A}}\cdot(\hat{\gamma_{j}}_{q}-Z_{\hat{\gamma_{j}}})((X_{q}(i,j)-Z_{X})+\hat{\beta^{*}_{j_{q}}})+Z_{A}$ (19)
@@ V-B Experiments Setup
Each model configuration underwent 50 training sessions consisting of 100 epochs, with early stopping implemented to prevent overfitting. We used a batch size of 256 and the Adam optimizer with standard parameters ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$, $\epsilon\!=\!10^{-9}$). The learning rate was initialized at 0.001, with a scheduler having a step size of 3 and a decay factor $\gamma$ of 0.5 for dynamic adjustment during training. These training sessions were conducted on an NVIDIA GeForce RTX 2080 SUPER GPU, utilizing CUDA 11.0 and PyTorch 3.11 within the Ubuntu 20.04.6 LTS. The objective metric for training was the minimization of the Mean Squared Error. The Mean Squared Loss function guided the training process. Post-training, we applied an inverse transformation to the model’s outputs and normalized target values to compute the Root Mean Square Error (RMSE) on the test data as the evaluation metric.
To generate the accelerator, Python scripts were used to translate the trained quantized model by passing its model and quantization parameters to VHDL templates, which resulted in corresponding VHDL files. We then used GHDL simulations to estimate the number of clock cycles required per inference. The accelerator design was synthesized using Vivado, generating comprehensive reports on resource utilization, timing, and power consumption. Finally, the accelerators were validated on real hardware to assess the effectiveness and efficiency of our FPGA implementation.
@@ V-C1 Parameter Count and FP32 Models
Overall, the optimal configuration for minimizing test RMSE across both datasets was $n\!=\!18$ and $d_{\text{model}}\!=\!32$. Notably, our FP32 model outperformed benchmarks reported in [18] for the PeMS dataset, demonstrating a 16.06% improvement. Similarly, compared to FP32 results on the AirU dataset documented by Becnel et al. [8] (where $n\!=\!24$ and $d_{\text{model}}\!=\!64$), our model achieved a 3.20% enhancement.
@@ V-C2 Quantization and Model Precision
The optimal configuration of the FP32 model did not consistently yield the best precision under quantization. For instance, on the AirU dataset, the optimal configuration (with $n\!=\!6$ and $d_{\text{model}}\!=\!64$) under 8-bit quantization resulted in 2.69% lower RMSE compared to the best FP32 model. Under 6-bit quantization, it (with $n\!=\!12$ and $d_{\text{model}}\!=\!64$) had 4.44% higher RMSE than the best FP32 model. For 4-bit quantization, $n\!=\!18$ and $d_{\text{model}}\!=\!64$ provided the best outcomes, with RMSE 46.71% higher than the best FP32 model. Notably, on the PeMS dataset, our 8-bit quantized model outperformed the 8-bit quantized model reported in [18] by 8.47%. Even our 6-bit quantized model delivered comparable performance. In addition, on the AirU dataset, our 8-bit quantized model surpassed the 8-bit quantized model reported in [8] by 33.47%. Remarkably, even our 4-bit quantized model outperformed their 8-bit counterpart by 2.83%.
@@ V-D Resource Utilization
Due to space constraints, this section focuses on the AirU dataset and evaluates resource utilization, including Lookup Tables (LUTs), Block RAM (BRAM), and Digital Signal Processing Slices (DSPs), for various model configurations and quantization bitwidths on the XC7S15 FPGA. As illustrated in Table III, even when quantized to 4-bit, the largest configuration ($n\!=\!24$ and $d_{\text{model}}\!=\!64$) exceeded the FPGA’s resource capacity, hence could not be deployed (denoted by “-”). Conversely, the smallest configuration ($n\!=\!6$ and $d_{\text{model}}\!=\!8$) fit easily within the FPGA’s constraints across different quantization bitwidths, with consistent BRAM utilization at 10%. LUTs and DSPs utilization increase with bitwidth, indicating higher logic resource demand with greater numeric precision.
@@
As discussed in Section V-C, the configuration with $n\!=\!6$ and $d_{\text{model}}\!=\!64$ under 8-bit quantization achieved optimal RMSE. According to Table III, this configuration barely fit the FPGA, utilizing at least 89.5% of all types of resources. For 6-bit quantization, the optimal RMSE was attained with the configuration of $n\!=\!12$ and $d_{\text{model}}\!=\!64$. Although its test RMSE is 7.33% higher than our best 8-bit quantized model, its LUTs utilization is 14.2% lower, while DSPs utilization increases by 5%. Among all 4-bit quantized models, the two most precise ones were too large to be deployed. We chose the model with the third-best RMSE performance (5.474), which has a 56.13% higher RMSE than our best 8-bit quantized model. However, its RMSE is only 0.63% higher than that of the 8-bit quantized model in [8]. This model, with the configuration of $n\!=\!12$ and $d_{\text{model}}\!=\!32$, consumes only 46.8% LUTs, 40% BRAM, and 65% DSPs. It represents a feasible compromise between resource utilization and precision, making it suitable for resource-constrained applications.
@@ V-E Timing Analysis
The fourth column of Table IV presents the number of clock cycles required per inference, which is influenced by model configuration ($n$, $d_{\text{model}}$). Comparing Rows 2 and 3, doubling $n$ resulted in 2.03$\times$ more clock cycles per inference. Similarly, comparing Rows 3 and 4, halving $d_{\text{model}}$ led to 3.46$\times$ fewer clock cycles per inference. The fifth column of Table IV outlines the model inference time, which depends on the clock frequency and the number of clock cycles. Notably, our 4-bit quantized model demonstrated the shortest inference time, while the 6-bit quantized model exhibited the longest.
@@ V-F Power and Energy Consumption
Energy consumption, as shown in the last column of Table IV, is calculated based on power usage and inference time. The 4-bit quantized model was the most energy-efficient, owing to its shorter inference time and lower power consumption. In contrast, the 6-bit quantized model underperformed the others, incurring the highest costs in both power and time. Notably, despite the 4-bit quantized model exhibiting a 56.13% higher RMSE than the 8-bit model, it is 2.12 times faster and 2.52$\times$ more energy-efficient. This efficiency highlights the potential advantages of deploying the 4-bit quantized model on smaller FPGAs, where resource constraints are more pronounced. Conversely, our 8-bit quantized model remains a suitable choice when precision is paramount.
@@ VI Related Work
Research on quantizing Transformer-based architectures has been extensive, particularly in the domains of NLP [13] and CV [14]. These models, known for their substantial computational demands, are primarily deployed on cloud-based servers and edge servers equipped with GPUs. However, relevant studies in TS, especially time-series forecasting, remain underexplored. Becnel et al. [8] implemented an 8-bit quantized Transformer on a low-power microcontroller unit (MCU) for time-series forecasting, achieving a power consumption of 23 mW but with an inference time of 176 ms. In contrast, our 4-bit quantized Transformer accelerator on an embedded FPGA achieves up to 132.33$\times$ faster inference with only a 0.63% increase in test loss. Despite a 2.74$\times$ increase in power consumption, it consumes 48.19$\times$ less energy, indicating that integrating an FPGA-based accelerator with the MCU could be beneficial.

### v5

@@ Abstract
[…]
@@ I Introduction
[…]
@@ IV-A Addition Component
$A^{3}=A^{1}+A^{2}\\$ (7)
$A^{3}\approx S_{A^{1}}\cdot(A^{1}_{q}-Z_{A^{1}})+S_{A^{2}}\cdot(A^{2}_{q}-Z_{A^{2}})\\$ (8)
$A^{3}_{q}\approx\left(\frac{S_{A^{1}}}{S_{A^{3}}}\cdot(A^{1}_{q}-Z_{A^{1}})+\frac{S_{A^{2}}}{S_{A^{3}}}\cdot(A^{2}_{q}-Z_{A^{2}})\right)+Z_{A^{3}}\\$ (9)
@@ IV-C Softmax Component
$S_{E}=\frac{1}{((2^{(2b-1)}-1)-(-2^{(2b-1)}))/(n^{2}\cdot h)}\\$ (12)
$Z_{E}=2^{(2b-1)}-\frac{1}{S_{E}}\\$ (13)
$\text{DLUT}(\hat{X}_{q})=\text{clamp}(\text{round}(\frac{E}{S_{E}}),-2^{2b-1},2^{2b-1}-1)\\$ (14)
$\text{NLUT}(\hat{X}_{q})=\text{clamp}(\text{round}(\frac{E}{S_{E}\cdot S_{A}}),-2^{3b-1},2^{3b-1}-1)\\$ (15)
@@ IV-D Batch Normalization Component
$A_{q}(i,j)$ $\approx\frac{S_{\hat{\gamma_{j}}}\,S_{X}}{S_{A}}\Bigl[\bigl(\hat{\gamma}_{j_{q}}-Z_{\hat{\gamma_{j}}}\bigr)\bigl(X_{q}(i,j)-Z_{X}\bigr)+\hat{\beta}_{j_{q}}^{*}\Bigr]+Z_{A}$ (19)
@@ V-B Experiments Setup
Each model configuration underwent 50 training sessions consisting of 100 epochs, with early stopping implemented to prevent overfitting. We used a batch size of 256 and the Adam optimizer with standard parameters ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$, $\epsilon\!=\!10^{-9}$). The learning rate was initialized at 0.001, with a scheduler having a step size of 3 and a decay factor $\gamma$ of 0.5 for dynamic adjustment during training. These training sessions were conducted on an NVIDIA GeForce RTX 2080 SUPER GPU, utilizing CUDA 11.0 and PyTorch 3.11 within the Ubuntu 20.04.6 LTS. The objective metric for training was the minimization of the Mean Squared Error. The Mean Squared Loss function guided the training process. Post-training, we applied an inverse transformation to the model’s outputs and normalized target values to compute the Root Mean Square Error (RMSE) on the test data as the evaluation metric. To generate the accelerator, Python scripts were used to translate the trained quantized model by passing its model and quantization parameters to VHDL templates, which resulted in corresponding VHDL files. We then used GHDL simulations to estimate the number of clock cycles required per inference. The accelerator design was synthesized using Vivado, generating comprehensive reports on resource utilization, timing, and power consumption. Finally, the accelerators were validated on real hardware to assess the effectiveness and efficiency of our FPGA implementation. Relevant source code is provided in the accompanying GitHub repository [^5] https://github.com/tianheng-ling/TinyTransformer4TS.
@@ V-C1 Parameter Count and FP32 Models
Overall, the optimal configuration for minimizing test RMSE across both datasets was n=18n\!=\!18 and $d_{\text{model}}\!=\!32$. Notably, our FP32 model outperformed benchmarks reported in [18] for the PeMS dataset, demonstrating a 16.06% improvement. Similarly, compared to FP32 results on the AirU dataset documented by Becnel et al. [8] (where n=24n\!=\!24 and $d_{\text{model}}\!=\!64$), our model achieved a 3.20% enhancement.
@@ V-C2 Quantization and Model Precision
The optimal configuration of the FP32 model did not consistently yield the best precision under quantization. For instance, on the AirU dataset, the optimal configuration (with n=6n\!=\!6 and $d_{\text{model}}\!=\!64$) under 8-bit quantization resulted in 2.69% lower RMSE compared to the best FP32 model. Under 6-bit quantization, it (with n=12n\!=\!12 and $d_{\text{model}}\!=\!64$) had 4.44% higher RMSE than the best FP32 model. For 4-bit quantization, n=18n\!=\!18 and $d_{\text{model}}\!=\!64$ provided the best outcomes, with RMSE 46.71% higher than the best FP32 model. Notably, on the PeMS dataset, our 8-bit quantized model outperformed the 8-bit quantized model reported in [18] by 8.47%. Even our 6-bit quantized model delivered comparable performance. In addition, on the AirU dataset, our 8-bit quantized model surpassed the 8-bit quantized model reported in [8] by 33.47%. Remarkably, even our 4-bit quantized model outperformed their 8-bit counterpart by 2.83%.
@@ V-D Resource Utilization
Due to space constraints, this section focuses on the AirU dataset and evaluates resource utilization, including Lookup Tables (LUTs), Block RAM (BRAM), and Digital Signal Processing Slices (DSPs), for various model configurations and quantization bitwidths on the XC7S15 FPGA. As illustrated in Table III, even when quantized to 4-bit, the largest configuration (n=24n\!=\!24 and $d_{\text{model}}\!=\!64$) exceeded the FPGA’s resource capacity, hence could not be deployed (denoted by “-”). Conversely, the smallest configuration (n=6n\!=\!6 and $d_{\text{model}}\!=\!8$) fit easily within the FPGA’s constraints across different quantization bitwidths, with consistent BRAM utilization at 10%. LUTs and DSPs utilization increase with bitwidth, indicating higher logic resource demand with greater numeric precision.
@@
As discussed in Section V-C, the configuration with n=6n\!=\!6 and $d_{\text{model}}\!=\!64$ under 8-bit quantization achieved optimal RMSE. According to Table III, this configuration barely fit the FPGA, utilizing at least 89.5% of all types of resources. For 6-bit quantization, the optimal RMSE was attained with the configuration of n=12n\!=\!12 and $d_{\text{model}}\!=\!64$. Although its test RMSE is 7.33% higher than our best 8-bit quantized model, its LUTs utilization is 14.2% lower, while DSPs utilization increases by 5%. Among all 4-bit quantized models, the two most precise ones were too large to be deployed. We chose the model with the third-best RMSE performance (5.474), which has a 56.13% higher RMSE than our best 8-bit quantized model. However, its RMSE is only 0.63% higher than that of the 8-bit quantized model in [8]. This model, with the configuration of n=12n\!=\!12 and $d_{\text{model}}\!=\!32$, consumes only 46.8% LUTs, 40% BRAM, and 65% DSPs. It represents a feasible compromise between resource utilization and precision, making it suitable for resource-constrained applications.
@@ V-E Timing Analysis
[…]
@@ V-F Power and Energy Consumption
[…]
@@ VI Related Work
[…]